Two cars leave at the same time traveling in different directions. One goes North at 6 mph and the other goes west at 8mph. How far apart are they after 4 hrs.
Answer by checkley77(12569)

I should have explained the method to solve this problem.
What you have is a right triangle which you know lengths of 2 of the sides which are the distances they travel.
These sides are:
6mph*4hrs=24 miles.
8mph*4hours=32 miles.
Now we have the formula for the hypotenuse (long side opposite the right angle) using the equation a^2+b^2=c^2.
I've substituted d for distance instead of the c value.
That's where I got the (^2) term.
---------------------------------------------------------------------
(6*4)^2+(8*4)^2=d^2
24^2+32*2=d^2
576+1024=d^2
1600=d^2
d=sqrt1600
d=40 miles is the distance between two cars after 4 hours